OpenStudy (anonymous):

Michelle draws a card from a standard deck of 52 cards. She replaces the card and draws a second card. What is the probability that she draws a red card and then a black card? A) 1/16 B) 1/26 C) 1/4 D) 1/52

Directrix (directrix):

Of the 52 cards in the deck, 26 are red and 26 are black. The probability of drawing a red card is 26/52. Whatever she draws, she places back in the deck. For the second draw, she is looking for a black card out of 52 cards in the deck. The probability of a black card is 26/52.

Directrix (directrix):

Putting the components of the problem together, P (R and B) = 26/52 * 26/52 = 1/2 * 1/2 = ? @breezymeetee

Directrix (directrix):

Answer is ....

OpenStudy (anonymous):

Multiply all the numbers?

OpenStudy (anonymous):

then simplify in lowest terms

Directrix (directrix):

1/2 * 1/2 = ?

OpenStudy (anonymous):

so it would be C?

Directrix (directrix):

I reduced the 26/52 to 1/2. If you have options that appear different, you might want to post them.

Directrix (directrix):

yes. C) 1/4

OpenStudy (anonymous):

The Options are: A) 1/16 B) 1/26 C) 1/4 D) 1/52

Directrix (directrix):

Okay. The answer is still c) 1/4

Directrix (directrix):

That comes from this: P (R and B) = 26/52 * 26/52 = 1/2 * 1/2 =

Directrix (directrix):

Question?

OpenStudy (anonymous):

what do u mean